Suppose you have $35,000 to invest. You’re considering Miller-Moore Equine Enterprises (MMEE), which is currently selling for $70 per share. You also notice that a call option with a $70 strike price and six months to maturity is available. The premium is $3.5. MMEE pays no dividends. What is your annualized return from these two investments if, in six months, MMEE is selling for $76 per share? What about $66 per share?

Annualized Return                         Stock           Option      
  $76 per share % %
  $66 per share % %  
Answer #1

NOTE: Assumption - One option buys only one unit of the underlying stock.

Stock:

Purchase Price of Stock = $ 70 and Price after 6 months = $ 76

6-Month Return = [(76-70) / 70] x 100 ~ 8.57 %

Annualized Return = 8.57 x 2 = 17.14 %

Purchase Price of Stock = $ 70 and Price after 6 months = $ 66

6- Month Return = [(66-70) / 70] x 100 ~ - 5,71 %

Annualized Return = -5.71 x 2 = - 11.42 %

Option:

Initial Investment = $ 35000 and Option Premium = $ 3.5,Number of Options Bought = 35000 / 3.5 = 10000

Stock Price after 6 months = $ 76

Option Payoff = (76 - 70) x 10000 = $ 60000

6-month return = [(60000/35000)-1] x 100 ~ 71.43 %

Annualized Return = 2 x 71.43 = 142.86 %

If the stock price is $ 66, the option position will have no payoff,thereby making the 6-month return [(0/35000) -1] x 100 = - 100 %

Annualized Return = -100 x 2 = -200 %